OpenStudy (anonymous):

Eliminate the parameter. x = t^2 + 1, y = t^2 - 1

OpenStudy (loser66):

from x = t^2 +1, t^2 =??

OpenStudy (anonymous):

-1

OpenStudy (loser66):

redo!! where is your x?

OpenStudy (anonymous):

x = t^2 + 1 t^2 = x-1 sorry lol I rushed a bit

OpenStudy (loser66):

same as y , from y =t^2 -1, t^2 =?

OpenStudy (anonymous):

t^2 = y+1

OpenStudy (loser66):

hence x-1 = y +1, or y = ??

OpenStudy (anonymous):

i believe its y = x-2 then

OpenStudy (loser66):

dat sit!! you got it

OpenStudy (anonymous):

awesome thanks, can you help with a few more?

OpenStudy (loser66):

not sure before seeing the problem.

OpenStudy (anonymous):

ok here: Eliminate the parameter. x = 5t, y = t + 8

OpenStudy (loser66):

do the same, isolate t for both x and y, let them equal and done.

OpenStudy (anonymous):

ok I am going to do it tell me if its right

OpenStudy (anonymous):

t = x/5 t = y-8 correct?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

then x/5 = y-8 y = x/5 + 8 is that it?

OpenStudy (loser66):

yup

OpenStudy (loser66):

easy, right?

OpenStudy (anonymous):

ok I have two more but I want to work them and then run them by if thats ok